Given that f is the real valued function defined by f(x) = 1 / (1 + e^x). And I = (1/4, 1/2).
(i)Show that if x E I, then f(x) E I. Where E represents "an element of".
(ii) Calculate f''(x), hence show that if x E I, thenㅣ f'(x)ㅣ<= 1/4.
(iii) Deduce that if x E I, then ㅣf(x) - sㅣ<= 1/4 ㅣx - sㅣ
s E (1/4, 1/2).
1
Expert's answer
2021-01-26T04:33:46-0500

(i) "x\in I, I=(\dfrac{1}{4}, \ \dfrac{1}{2})"


"f(x)=\dfrac{1}{1+e^x}"

"f'(x)=(\dfrac{1}{1+e^x})'=-\dfrac{e^x}{(1+e^x)^2}<0, x\in \R"

Therefore the function "f(x)" decreases for "x\in \R."


"f(\dfrac{1}{4})=\dfrac{1}{1+e^{1/4}}<\dfrac{1}{1+(1)^{1/4}}=\dfrac{1}{1+1}=\dfrac{1}{2}"

"f(\dfrac{1}{2})=\dfrac{1}{1+e^{1/2}}>\dfrac{1}{1+(4)^{1/2}}=\dfrac{1}{1+2}=\dfrac{1}{3}>\dfrac{1}{4}"

Therefore


"\dfrac{1}{4}<f(x)<\dfrac{1}{2}, x\in(\dfrac{1}{4},\ \dfrac{1}{2})"

Therefore if "x\in I," then "f(x)\in I."


(ii)


"f'(x)=(\dfrac{1}{1+e^x})'=-\dfrac{e^x}{(1+e^x)^2}<0, x\in \R"

"f''(x)=(-\dfrac{e^x}{(1+e^x)^2})'"

"=-\dfrac{e^x(1+e^x)^2-2e^{2x}(1+e^x)}{(1+e^x)^4}"

"=\dfrac{e^x(e^x-1)}{(1+e^x)^3}>0, x>0"

Therefore for "x>0" the function "f'(x)" increases and "f(x)<0."


Therefore for "x>0" the function "|f'(x) |" decreases and "|f(x)|>0."


"|f'(x)|<|f'(0)|=\dfrac{e^{0}}{(1+e^{0})^2}=\dfrac{1}{4}, x\in I"

"|f'(x)|\leq\dfrac{1}{4}, x\in I"

(iii)

Let "x=0.3, s=0.3"


"f(0.3)=\dfrac{1}{1+e^{0.3}}>\dfrac{1}{1+1.5}=0.4"


"|f(x)-s|>|0.4-0.3|=0.1"

"\dfrac{1}{4}|x-s|=\dfrac{1}{4}|0.3-0.3|=0"

Hence the statement


"|f(x)-s|\leq\dfrac{1}{4}|x-s|, s\in(\dfrac{1}{4}, \ \dfrac{1}{2})"

is False.